How can you solve this with elimination?
3x + 2y= -6
-x - 2y= 10

Answer by checkley79(3341)

3x + 2y= -6
-x - 2y= 10 ADD.
----------------------
2X=4
X=4/2
X=2 ANS.
-2-2Y=10
-2Y=10+2
-2Y=12
Y=12/-2
Y=-6 ANS.
PROOF:
-2-2(-6)=10
-2+12=10
10=10